HUBBLE SPACE TELESCOPE DAILY REPORT #904 PERIOD COVERED: 0800 EST Daily Status Report as of 160/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOC/96 4353 (Hubble space Telescope: Imaging of Io: Cycle 3 High) The Faint Object Camera was used to make one ultraviolet observation of Jupiter's moon Io in order to determine the abundance and distribution of surface and atmospheric constituents, particularly sulfur dioxide. The first observations was a dark calibration exposure taken off target to check out the camera. The OSS then sent a corrective maneuver to the spacecraft to move the aperture to the target and to verify the filters. The image of the target was displayed in the OSS, and was reported to look normal. 1.2 Completed Three Sets of WFC 4713 (WF/PC Cycle Two Calibration: Internal Delta Flats) A total of seven internal flat calibration images were made using the Wide Field Camera. These images will be used for the restoration of Earth flats, which were altered due to changes in quantum efficiency which occurred after WFPC decontamination operations. All seven observations were monitored in the OSS. All were reported to be normal. 1.3 Completed FOS/RD 4772 (Black Holes, Stellar Dynamics and Populations in Nuclei) The Faint Object Spectrograph was used to study the central core nucleus region of elliptical galaxy NGC-4473 to study its stellar population and dynamics of its core. The peakups appeared to be successful. However, the conformation image was reviewed and appeared to be blank. 1.4 Completed WFC 4381 (HST Medium-Deep Survey - Cycle 3) The Wide Field Camera was used to make one medium-deep high latitude sky survey image taken in parallel with the prime science observation. The observation was monitored and displayed. It was reported to be normal. 1.5 Completed WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) Both the Wide Field Camera and the Planetary Camera were used to make on internal flat observation each. These observations were to check for new forms of contamination (i.e., daisies and measles), and the possible return of quantum efficiency hysteresis which occurs after decontamination operations. Both images were received and displayed in the OSS. They were reported to be normal. 1.6 Completed PC 4140 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor - Part 2) The Planetary Camera was used to make two two rapid internal monitor observations (internal flats) to check for contamination. Both observations occurred during LOS. They were later displayed and reviewed. The OSS reported that they were normal. 1.7 Completed WFC 4929 (WF/PC Cycle 2 : Improved Dark Calibration) The Wide Field Camera was used to make a dark calibrations image. The Observation lasted 1200 seconds, and the camera was pre-flashed. The observations were executed as scheduled, and no problems were reported. 1.8 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.9 Completed FOC/96 3804 (Imaging of UV Bright Stars In Metal Rich Globular Clusters) The Faint Object Camera (f/96) was used to make far-ultrviolet observations of metal-rich stars in the globular cluster NGC-5921 to determine their temperatures and luminosity. The observations were monitored by the OSS, and the data quality was assessed. No problems were reported. 1.10 Completed PC 4902 (HST Medium-Deep Survey - Part 2: Low-Lat - Cycle 3 High) The Planetary Camera was used to make some low latitude medium-deep sky survey images taken in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. 1.11 Completed FOC/96 4804 (UV Imaging of Nearby Galaxies FOC/96) The Faint Object Camera (f/96) was used to make ultraviolet snapshot images of nearby galaxy ESO-1335-1 looking for a possible active galactic nucleus (aka, AGN). The exposure lasted ten minutes and used filter F220W. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 5 Successful : 5 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Significant HSTARs: None 2.4 Operations Note: Flight Software --------------- An installation of version 9.0 of the DF-224 flight software was initiated as planned this morning (June 9, 160/1200Z). This will be the first time a complete new DF-224 version has been installed on the spacecraft since launch. A successful pre-Flight Readiness Review for this installation was held this week. No problems have been reported. If all goes as planned the SMS-161 will be initiated at 161/0800Z. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete S/C 4974 (The Installation of Flight Software 9.0) 3.2 Complete S/C VSAFE (Safemode recovery of the HST Science Instruments - HSP ) 3.3 Complete S/C XSAFE (Safemode recovery of the HST Science Instruments - FOC ) 3.4 Complete S/C YSAFE (Safemode recovery of the HST Science Instruments - FOS ) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-